Mall security estimates that the average daily per-store theft is less than$250, but wants to determine the accuracy of this statistic. The company research takes a sample of 81 clerks and finds that X=$260 and s = $20
a) Test x =.05
B) CONSTRUCT A 95% cie of u

Answer by stanbon(75887)

Mall security estimates that the average daily per-store theft is less than$250, but wants to determine the accuracy of this statistic. The company research takes a sample of 81 clerks and finds that X=$260 and s = $20
a) Test x =.05
Ho: u = 250
Ha: u < 250 (claim)
---
z(260) = (260-250)/[20/sqrt(81)] = 10/(20/9) = 90/20 = 4.5
p-value = P(z < 4.5) = 0.9999
Since the p-value is greater than 5%, fail to reject Ho
The test results do not support the claim.
=============================================
B) CONSTRUCT A 95% cie of u
x-bar = 260
ME = 1.96 = 4.36
----
95% CI: 260-4.36 < u < 260+4.36
(255.64,264.36)
=====================
Cheers,
Stan H.
======================